What are all the possible sets of three consecutive odd whole numbers whose sum is less then 45?
Answer by oberobic(2304)

Three consecutive odd whole numbers are x, y, and z.
But with three unknowns, we need three equations.
All we have is 1 equation, so we need to define the numbers differently.
x
x+2
x+4
.
x +x+2 +x+4 = 45
.
3x +6 = 45
.
3x = 39
.
x = 13
x+2 = 15
x+4 = 17
.
13+15+17=45
.
Answer: The three consecutive whole odd numbers that total 45 are 13, 15, and 17.
.
Done.